OpenStudy (dls):

If(a1,a2),(g1,g2),(h1,h2) are respectively pairs of AM's GM's and HM's inserted between two numbers then find h1+h2/h1h2

OpenStudy (anonymous):

there is a relation connecting these three u knw that

OpenStudy (dls):

I guess no :/

OpenStudy (anonymous):

G^2 = A * H

OpenStudy (dls):

thanks for this !

OpenStudy (anonymous):

u may ..use this...)) here

OpenStudy (dls):

would it be something like h1h2=g1g2/a1a2?

OpenStudy (anonymous):

g1^2 = a1 * h1 g2^2 = a2 * h2

OpenStudy (dls):

oh :o

OpenStudy (anonymous):

Wat..))

OpenStudy (dls):

what after this

OpenStudy (dls):

do i have to multiply both the equations?

OpenStudy (anonymous):

find h1 and h2 from this......and eq to h1+h2/h1h2

OpenStudy (dls):

g1^2 g2^2 / a1a2 = h1h2

OpenStudy (dls):